A bird is flying with a speed of 16.8 m/s over water when it accidentally drops a 2.50 kg fish.

The acceleration of gravity is 9.81 m/s2 .

If the altitude of the bird is 6.20 m and air resistance is disregarded, what is the speed of the fish when it hits the water?

Answer in units of m/s.


1
Expert's answer
2021-12-13T11:08:06-0500

"V=16.8m/sec\\m=2.50kg\\g=9.81m/sec^2\\h=6.20m"


"Machanical energy ( ME)=\frac{1}{2}mv^2+mgh"

"ME=\frac{1}{2}\times2.50\times16.8^2+2.50\times9.81\times6.20"

ME =352.8+152.055=504.855J

E=504.855J

"V=\sqrt {\frac{2E}{m}}"

"V=\sqrt{\frac{2\times504.855}{2.50}}"

"V=20.09m\/sec"